OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Solve the following equation for w: P = 2l + 2w a. w=p/2-1 b. w=p/2+1 c. w=p-1/2 d. w=p+1/2

OpenStudy (anonymous):

is it 21 or 2l ?

OpenStudy (anonymous):

2l it looks like

OpenStudy (anonymous):

To solve for W, we need to isolate it to one side of the equation -- doesn't matter which side. . P = 2W + 2L Start by subtracting 2L from both sides:

OpenStudy (anonymous):

P - 2L = 2W Finally, divide both sides by 2: (P-2L)/2 = W or, you could write: P/2 - L = W

OpenStudy (anonymous):

therefore the answer is A